Store A sells jeans for 7/8 of the price at store B. Store A sells jeans for $35. Write and solve an equation to find out how much you save by buying jeans at store A.
Answer by rfer(16322)

35/(7/8)=
35*8/7=40 store B
40-35=$5 Saved